OpenStudy (anonymous):

What is the value of the expression when a = 4, b = 8, and c = 5?

OpenStudy (anonymous):

this is the question

OpenStudy (igreen):

What's the expression? @jpharper

OpenStudy (anonymous):

right ubove

OpenStudy (eric_d):

Just plig in the given values

OpenStudy (eric_d):

pug*

OpenStudy (eric_d):

into the above expression

OpenStudy (eric_d):

You may try that now @jpharper

OpenStudy (eric_d):



OpenStudy (igreen):

a = 4, b = 8, and c = 5 \(\dfrac{3a + b}{2c}\) \(\dfrac{3(4) + 8}{2(5)}\) \(\dfrac{12 + 8}{10}\) \(\dfrac{20}{10}\) \(2\)

OpenStudy (anonymous):

cool thanks